OpenStudy (anonymous):

How do you simplify 10 over radical seven plus radical two?

OpenStudy (anonymous):

\[(10)/(\sqrt{2} +\sqrt{2})\]?

OpenStudy (radar):

\[10/\sqrt{7+\sqrt{2}}\]

OpenStudy (anonymous):

It is \[5/\sqrt{2}\]

OpenStudy (anonymous):

\[\sqrt{2} + \sqrt{2} = 2\sqrt{2}\] Therefore 10/ 2sqrt(2) will equal what I have above

OpenStudy (radar):

What happened to the 7 that was under the radical?

OpenStudy (radar):

rpeterson is the problem

OpenStudy (anonymous):

What was written was different than if it is 7 instead of 2 you are correct but if its 2 than I am correct

OpenStudy (radar):

\[10/(\sqrt{7}+\sqrt{2}\]

OpenStudy (radar):

rpeterson there seems to be some confusion on your problem. Please restate the problem

OpenStudy (anonymous):

Where is he?

OpenStudy (radar):

I don't know at first i thought there was two radicals one inside the other as I posted, but now I think there is two separate radicals.

OpenStudy (anonymous):

\[(10)\div \sqrt{7}+\sqrt{2}\] ? or : \[(10)/\sqrt{7+\sqrt{2}}\]

OpenStudy (anonymous):

I think just wait he come back

OpenStudy (radar):

If it is \[(10)/(\sqrt{7}+\sqrt{2})\] then\[2(\sqrt{7}-\sqrt{2})\]

OpenStudy (anonymous):

yes, right

OpenStudy (anonymous):

how about:\[(10)/\sqrt{7+\sqrt{2}}\]

OpenStudy (radar):

If it is like that Nancy Lam, I don't want to attempt to solve it lol